### v3

# Autoregressive Stochastic Clock Jitter Compensation in Analog-to-Digital Converters

Abstract
This paper deals with the mathematical modeling and compensation of stochastic discrete time clock jitter in Analog-to-Digital Converters (ADCs). Two novel, computationally efficient de‑jittering sample pilots‑based algorithms for baseband signals are proposed: one consisting in solving a sequence of weighted least‑squares problems and another that fully leverages the correlated jitter structure in a Kalman filter‑type routine. Alongside, a comprehensive and rigorous mathematical analysis of the linearization errors committed is presented, and the work is complemented with extensive synthetic simulations and performance benchmarking with the scope of gauging and stress‑testing the techniques in different scenarios.
Index Terms:
Analog-to-Digital Converters, stochastic jitter, autoregressive process, weighted least-squares, Kalman smoother, linearization.
I Introduction
Analog-to-digital converters introduce various distortions into sampled signals. These arise from hardware imperfections and appear as DC offset biases, nonlinear distortion from clipping, and irregularities in sampling instants (jitter) [43].
The ADC clock typically uses a voltage controlled oscillator locked to a reference clock by a phase locked loop with a defined loop bandwidth. Inside that bandwidth, loop noise relates in scale to the reference phase noise; outside it, noise originates from the oscillator itself and the clock distribution network. When converted to time, this phase noise appears as clock jitter. Cumulative phase fluctuations from both the loop regulated portion and the free running oscillator directly add to time jitter, degrading the timing accuracy and overall performance of the converter.
In the absence of noise, perfect recovery of a deterministic but non uniform sampling pattern can be achieved using classical Paley-Wiener-Levinson theory [25]. Driven by advances in converter technology, work on random timing offsets began in the early 1960s [8][4]. In that setting optimality is measured not at each sample point but in terms of mean squared error [44][32][15].
Modern work on identifying and compensating jitter in converters has included techniques based on injecting pilot tones [40][34][42]. When pilot sinusoids are present, jitter appears as phase noise disturbance, allowing dedicated phase noise tracking and compensation methods to be applied [40][34][42]. Although jitter is often modeled as uncorrelated additive Gaussian uncertainty [44][41][3][20], the clock jitter process neither has a white spectrum nor is generally uncorrelated [9]. Conflicting models of its power spectral density appear in the literature - for example [42] versus [11][46] -particularly in their treatment near zero Hertz. However, all surveyed models agree on a $1/f^{2}$ spectral decay in mid-band decades. That property, along with correlated behavior capturing slow and fast jitter [9], can be modeled by autoregressive (AR) processes or more generally by autoregressive moving-average (ARMA) processes. Experiments reported in [46] show a narrow region near DC exhibiting a $1/f^{3}$ decay. Capturing that requires more flexible models such as autoregressive fractionally integrated moving average (ARFIMA) processes, which introduce additional complications from potential non-stationarity. Standard Kalman filtering must also be adapted to handle state-space models with fractional operators [39].
The main contribution of this paper is summarized as follows:
[…]
    Upon modeling the stochastic clock jitter as an autoregressive process of order $1$ with parameter close to $1$, which is simple, mathematically tractable and mid-band “spectrally faithful”, we propose and compare two novel jitter tracking and compensation algorithms based on pilot samples; one employs Kalman filtering and smoothing to fully exploit the jitter mathematical structure, the other leverages optimal weighted least-squares and it is more “model agnostic” (Section III).
Furthermore the analysis is complemented and enriched with:
[…]
    rigorous justifications for the approximations arising in Taylor expansions and outline of the regime in which these approximations are valid (Section II-C);
[…]
    synthetic simulations to stress test the techniques presented (Section IV).
I-A Notation and symbols
ADCs act on continuous-time signals by discretizing and digitalizing them, so that the output of an ADC is a discrete sequence of values $x(t_{n})$ when the input is the analog $x(t)$. If $x(t)$ is a stochastic process, we can model the ADC output as a sequence of random variables; we will indicate sequences (finite or infinite) with bold lowercase letters $\mathbf{x}\coloneqq(x_{1},x_{2},\dots)$; if $\mathbf{x}$ and $\mathbf{y}$ are two such sequences, $\mathbf{x}\odot\mathbf{y}$ indicates their Hadamard componentwise product. Bold uppercase letters $\mathbf{A},\mathbf{B},\dots$ will be reserved to matrices; the symbol $\approx$” stands for “approximately equal to” (with some negligible error) while the symbol $\ll$” means “much less than”. With $\widehat{\cdot}$ we will indicate estimations of sought quantities, and with $\widetilde{\cdot}$ noisy measurements of quantities.
[…]
Derivatives
by $x^{\prime}_{n}=x^{\prime}(nT_{s})$ we mean the first-order time derivative of the process $x(t)$ calculated at times $nT_{s}$; for a bandlimited Gaussian process this operation is well-defined [5] and it is equivalent to the $n$-th sample of the discrete derivative operator $D$ applied to the (infinite) sequence of samples, i.e.
$\frac{d}{dt}x(t)\big{|}_{t=nT_{s}}=x^{\prime}(nT_{s})=(D\mathbf{x})_{n}.$
This derivative operator $D$ acts as time domain convolution
$(D\mathbf{x})_{n}=\frac{1}{T_{s}}\sum_{k}h_{k}x_{n-k}$
with the ideal non-casual bandlimited derivative filter $h_{k}=(-1)^{k}/k$ for $k\neq 0$ and $h_{0}=0$; its frequency response is $H_{D}(\Omega)=i\Omega$. These expressions is commonly found and used in the relevant literature, cfr. [35][13][26].
Oftentimes throughout the text we will encounter continuous-time processes $y$ (Wiener, Ornstein-Uhlenbeck, white noise) whose derivative might not exist anywhere in the standard pointwise sense. However, upon bandlimiting and sampling them, their discrete $D\mathbf{y}$ counterparts will exist; to keep the notation as light as possible we will write, with a slight abuse of notation, $y^{\prime}_{n}$ in lieu of $(D\mathbf{y})_{n}$.
II Signal and system model
We consider baseband analog signals that are modeled as continuous-time bandlimited complex-valued stationary Gaussian processes $x(\omega,t)$ so that, for each time instant $t\in T$, the random variable $(\omega\mapsto x(\omega,t))\sim\mathcal{N}(0,1)$; $x$ has flat power spectral density $\mathcal{S}_{x}(f)=1/(2W)\chi_{\{|f|\leq W\}}(f)$, where $\chi(\cdot)$ denotes the indicator function. This is a standard model used to describe signals transmitted in telecommunication applications, cfr. [30]. We recall that sample paths of these processes are analytically well-behaving, i.e. the maps $t\mapsto x(\omega,t)$ are entire (and thus differentiable with continuity) for almost every $\omega\in\Omega$, as direct consequence of e.g. Theorem 11 in [5]. Therefore the “n-th derivatives objects” $x^{(n)}$ are well-defined for all $n\in\mathbb{N}$.
II-A Stochastic jitter description and problem statement
In an ideal scenario, the ADC would sample signals at uniformly spaced intervals, meaning $t_{n+1}-t_{n}$ is constant for all $n$. However, imperfections in the local oscillator (LO) circuitry introduce jitter, causing the actual sampling instants to deviate from the ideal in a non-uniform way: i.e. the actual sampling is done at times $\tilde{t}_{n}=nT_{s}+\xi_{n}$ where $T_{s}$ is the ADC sampling interval and $\xi_{n}$ is the time-dependent (stochastic) jitter. The literature presents different models for the (discrete) process $\{\xi_{n}\}_{n\geq 1}$: in [32] $\xi_{n}\sim\mathcal{U}([-a,a])$ and independent while in [44][41][3] the $\xi_{n}$ are modeled as i.i.d. Gaussian with 00 mean and variance $\sigma^{2}$. [42] describes $\xi_{n}$ as a “slowly varying” Gaussian process. To the best of our knowledge only [9] models the time domain clock jitter as autoregressive process, while the Wiener process is more broadly used e.g. in [29][16][23]. Several sources focus on the spectral properties of this process and derive descriptions thereof; it turns out that those very formulas equal those expressing the power spectral densities of AR(1) processes; compare e.g. Formula (9) in [9] with Formula (2), or Formula (4) in [11] with Formula (3).
Therefore, in the present we model the jitter process as a discrete autoregressive process of order $1$ AR(1) with evolution relation
$\xi_{n}=\varphi\xi_{n-1}+\epsilon_{n},$ (1)
where $\varphi\approx 1$, $\epsilon_{n}\sim\mathcal{N}(0,\sigma_{\epsilon}^{2})$ and $\xi_{0}\sim\mathcal{N}(0,\sigma_{\epsilon}^{2}/(1-\varphi^{2}))$ (stationary initialization); it can be shown [22] that the power spectral density of this process is
$\mathcal{S}_{\xi}(e^{i\omega})=\frac{\sigma_{\epsilon}^{2}}{1+\varphi^{2}-2\varphi\cos(\omega)},\ \omega\in[-\pi,\pi],$ (2)
from which it is pretty straightforward to observe that much of the process power is concentrated near $\omega=0$ (or 00 Hz): indeed
$\begin{split}\mathcal{S}_{\xi}(e^{i\omega})&=\frac{\sigma_{\epsilon}^{2}}{(1-\varphi)^{2}+2\varphi(1-\cos(\omega))}\\ &=\frac{\sigma_{\epsilon}^{2}}{(1-\varphi)^{2}+4\varphi\sin^{2}(\omega/2)}\\ &\leq\frac{\sigma_{\epsilon}^{2}}{(1-\varphi)^{2}+4\varphi\omega^{2}/\pi^{2}}\end{split}$ (3)
where we used $\sin(\omega)\geq 2\omega/\pi$ for $\omega\in[0,\pi/2]$, from which it is again evident the desired quadratic decay in $\omega$. In fact, the last formula in (3) is the PSD of the Ornstein-Uhlenbeck process, that might be regarded as the continuous counterpart of an AR(1). The AR(1) jitter description formulation offers a faithful, convenient and mathematically tractable representation of the disturbance process, which appears to have been overlooked in the literature.
It has been observed [42] that $\xi_{n}$ could be weakly correlated with the white background noise, but in the present we will consider them as uncorrelated.
If the jitter is not too large (with respect to the sampling interval), its effect on the discretized signal is often [42][41][32][14] approximated by means of first-order Taylor expansion:
$x(nT_{s}+\xi_{n})\approx x(nT_{s})+\xi_{n}x^{\prime}(nT_{s}).$ (4)
As we show later in Proposition II.1, the higher order terms in the Taylor expansion can be considered in some sense negligible.
From here on, we will also set
$y_{n}\coloneqq x(nT_{s})+\xi_{n}x^{\prime}(nT_{s})+w_{n}$ (5)
as our model under investigation, with $w_{n}\sim\mathcal{N}(0,\sigma_{w}^{2})$ a sequence of independent Gaussian random variables modeling additional white noise in measurements, which are also independent of the process $\{\xi_{n}\}_{n\in\mathbb{N}}$.
The problem we aim to tackle in the present is to obtain an estimation of the process $\xi_{n}$ given measurements $y_{n}\ \forall n$ together with pilot samples $x_{n}$ for $n\in I\subseteq\mathbb{N}$ and, as byproduct, to derive a de-jittered estimation of $x_{n}$.
II-B How much jitter is “too much” jitter?
How much jitter is “too little” or “too much” highly depends on the specific application, but we still want to give a few raw numbers present in literature. From a theoretical perspective, [32] studies scenarios in which the ratio between signal bandwidth and jitter is up to $10\%$, and the ratio between sampling interval and jitter is up to $1\%$. Similarly, [42] discusses an algorithm for systems where the jitter is $0.5-1\%$ of the sampling rate. However lower values have been practically achieved already since the end of the 80s across multiple sampling rates ranges, e.g. in [17] ($5.5$ ps jitter at $1$ GS/s, $\sim 0.5\%$), [38] ($0.53$ ps jitter at $65$ MS/s, $\sim 0.0035\%$); Ali et al. [2] achieved $88$ fs at $125$ MS/s ($\sim 0.0011\%$).
One way to quantify how much jitter shall be tolerated is by setting a target Signal-to-Noise-and-Distortion Ratio (SINADR) value; since the disturbance is cast as additive (cfr. (4)), we can write the SINADR for our signal and jitter model (excluding quantization noise)
$\begin{split}\text{SINADR}_{\text{dB}}&=10\log_{10}\left[\frac{\text{var}(x_{n})}{\text{var}(w_{n})+\text{var}(\xi_{n}x^{\prime}_{n})}\right]\\ &=10\log_{10}\left[\frac{\text{var}(x_{n})}{\text{var}(w_{n})+\text{var}(\xi_{n})\text{var}(x^{\prime}_{n})}\right]\\ &=10\log_{10}\left[\frac{3(1-\varphi^{2})\sigma_{x}^{2}}{3(1-\varphi^{2})\sigma_{w}^{2}+4\pi^{2}W^{2}\sigma_{\epsilon}^{2}\sigma_{x}^{2}}\right],\end{split}$ (6)
where we used $\sigma_{\xi}^{2}=\sigma_{\epsilon}^{2}/(1-\varphi^{2})$, $\mathbb{E}[\xi_{n}]=\mathbb{E}[x^{\prime}_{n}]=0$ and the fact that $\xi_{n}$ and $x_{n}$ are independent. By means of inverting eq. (6) it is easy to see how much jitter could be tolerated (for a desired SINADR level). At the same time for the first-order Taylor expansion (4) to hold, the jitter $\xi_{n}$ cannot be too large; therefore one reasonable assumption is to make the “small jitter” hypothesis (SJH), which is in line with the literature surveyed so far, i.e. to request
$\sigma_{\xi}/T_{s}\ll 1;$
we will refer to the left hand side of the latter as jitter percentage. Notice that when $\sigma_{w}=0$, (6) becomes equal to
$-20\log_{10}\left[2\pi W\sigma_{\xi}/\sqrt{3}\right]\approx-20\log_{10}\left[2\pi W\sigma_{\xi}\right]+4.77\text{ dB},$
which is the standard jitter SNR formula found in classical literature (see e.g. [36]), applied to bandlimited signals.
The interaction between the two additive distortions in (5) will also play a role in our analysis; therefore we introduce a convenient notion of Noise-to-Distortion Ratio (NDR), an obvious variant of (6):
$\text{NDR}_{\text{dB}}=10\log_{10}\left[\frac{\sigma_{w}^{2}(1-\varphi^{2})}{4\pi^{2}W^{2}\sigma_{x}^{2}\sigma_{\epsilon}^{2}}\right].$ (6)
II-C Approximation errors analysis
There are two main approximations that will be performed in our analysis, whose errors need to be rigorously analyzed. In the first-order expansion in (4) under (II-B) the neglected higher-order terms are, in some sense, negligible, as the following Proposition clarifies:
Proposition II.1.
Let $\mathbf{x}=\{x_{n}\}_{n\geq 1}$ be a discrete zero-mean, wide-sense Gaussian stationary bandlimited process and $\boldsymbol{\xi}=\{\xi_{n}\}_{n\geq 1}$ a discrete autoregressive process of order $1$ with variance $\sigma_{\xi_{n}}^{2}\equiv\sigma_{\xi}^{2}$, independent on $\mathbf{x}$. Assume that (II-B) holds; then we have
$\emph{var}(\xi_{n}^{2}x^{\prime\prime}_{n})\ll\emph{var}(\xi_{n}x^{\prime}_{n})\quad\forall\,n.$ (7)
The second error source is in the derivative of $y$ (5), that will replace the unknown $x^{\prime}_{n}$ in our algorithms:
$y_{n}^{\prime}=x_{n}^{\prime}+u^{\prime}_{n}+w^{\prime}_{n},$ (8)
where we set $\mathbf{u}=\boldsymbol{\xi}\odot D\mathbf{x}$. The term $w^{\prime}_{n}$ is violet noise (bandlimited to $1/2T_{s}$), whose variance equals $\pi^{2}\sigma_{w}^{2}/3T_{s}^{2}$. For a fixed NDR level $r$, the white noise variance satisfies $\sigma_{w}^{2}=10^{r/10}\text{var}(\xi x^{\prime})$. Therefore
$\sigma_{w^{\prime}}^{2}=\frac{\pi^{2}10^{r/10}\sigma_{\xi}^{2}}{3T_{s}^{2}}\sigma_{x^{\prime}}^{2}\approx 3.29\frac{\sigma_{\xi}^{2}}{T_{s}^{2}}10^{r/10}\sigma_{x^{\prime}}^{2}$ (9)
and thus $\sigma_{w^{\prime}}^{2}\ll\sigma_{x^{\prime}}^{2}$ provided that (II-B) holds and the white noise is not too strong. The term $D(\boldsymbol{\xi}\odot\mathbf{x}^{\prime})_{n}$ has a more convoluted structure, but its impact can be explicitly assessed, as shown in Proposition II.2, whose proof can be found in the appendix. Figure 1 shows again that $\sigma_{x^{\prime}}^{2}$ is fully dominating.
Proposition II.2.
Let $\mathbf{x}=\{x_{n}\}_{n\geq 1}$ be a discrete zero-mean, wide-sense Gaussian stationary bandlimited process (sampled every $T_{s}$ seconds), with its derivative $\mathbf{x}^{\prime}=\{x_{n}^{\prime}\}_{n\geq 1}$, and $\boldsymbol{\xi}=\{\xi_{n}\}_{n\geq 1}$ a discrete autoregressive process of order $1$ with parameter $\varphi\in(0,1)$ and variance $\sigma_{\xi_{n}}^{2}\equiv\sigma_{\xi}^{2}$, independent on $\mathbf{x}$. Let also $D$ be the time-domain ideal derivative operator. Then the following holds:
$\emph{var}(D(\boldsymbol{\xi}\odot\mathbf{x}^{\prime})_{n})=\frac{\sigma_{\xi}^{2}}{T_{s}^{2}}\frac{3(1-\varphi^{2})}{32\pi^{4}(WT_{s})^{3}}I\sigma_{x^{\prime}}^{2}$ (10)
[…]
Remark II.1.
We point out that a stochastic version of the Peano remainder theorem for Taylor expansions hold, cfr. Lemma 3 in [45]. Further results comparing (conditional) variances of the objects $x(\omega,nT_{s}+\xi_{n})$ and $x(\omega,nT_{s})+x^{\prime}(\omega,nT_{s})\xi_{n}$ would require some care; first of all the measurability/randomness of the maps $\omega\mapsto x(\omega,nT_{s}+\xi_{n}(\omega))$ or $(\omega_{1},\omega_{2})\mapsto x(\omega_{1},nT_{s}+\xi_{n}(\omega_{2}))$ must be carefully assessed, depending on whether the jitter process is seen as living in the same probability space. Measurability of $x(\omega,t)$ on the whole product space $\Omega\times\mathbb{R}$ typically require left- or right-continuity in $t$ to hold for all $\omega\in\Omega$ (see Lemma 6.4.6 in [7]), which might not be the case for our signal model. In that case one shall replace $x$ with an indistinguishable copy for which the continuity holds for every $\omega\in\Omega$.
Figure 1: Level curves of the term multiplying $\sigma_{x^{\prime}}^{2}$ in (10); $\varphi=0.95$ left, $\varphi=0.999$ right, in dB. The double integral in (10) was numerically evaluated via nested integral in MATLAB and its accuracy was double-checked via a separate Monte Carlo where instances of the processes $\mathbf{x}$, $\boldsymbol{\xi}$ etc. were generated and the left-hand side of (10) estimated.
III Pilot samples-based jitter compensation
In this section are found the main contributions of this paper. We present two jitter tracking techniques aimed at mitigating the negative impact of this type of disturbance. Both algorithms will assume the knowledge of pilot samples at known time index positions, meaning that the true sample values at $n\in I\subsetneq\mathbb{N}$ will be assumed to be known a priori. This may achieved be programming the analog front end so that at specific time instants the ADC is fed with a known analog stimulus; the outputted stream of samples will be jittered, including the samples at pilot positions. A noisy estimation of the jitter process sample values can be obtained by reversing equations (4) and (5) at pilot positions:
$\xi_{n}\approx\widetilde{\xi}_{n}=(y_{n}-x_{n})/x^{\prime}_{n},\quad n\in I.$ (11)
Once the jitter process is estimated using one of the methods detailed in this section, the jitter distortion in (5) must be subtracted, the pilots removed and the signal interpolated using e.g. the Gerchberg-Papoulis algorithm [27][18] to “fill the holes” left after the removal of pilot samples. The workflow is summarized by the scheme in Figure 2; the focus of this paper is within the green box.
Introduce known analog stimulus. Use pilot samples to estimate $\xi_{n}\,\forall\,n$ and compensate the distorted signal. Remove pilot samples and interpolate (w e.g. Gerchberg-Papoulis).
Figure 2: Schematic three‐step process.
As opposed to pilot tones-based techniques [42], we shall note that no knowledge of the signal’s derivative is assumed by our techniques, i.e. $x^{\prime}_{n}$ is not assumed to be a priori known at pilot positions (nor anywhere else). Moreover, the knowledge of the signal’s samples $x_{n}$ at pilot positions is not enough to compute its derivative at these positions. In our algorithms we will use $y^{\prime}_{n}$ instead, whose utilization is justified by the analysis carried out in Section II-C. An example of tracking is showcased in Figure 3; despite the large amount of missing data, both methods are able to reconstruct the aggregated behaviour of the underlying autoregressive process.
Figure 3: Jitter trackers comparison example, Algorithm 1 (left) vs Algorithm 2 (right).
III-A Robust blockwise polynomial interpolation
The hidden true process is treated here as blockwise deterministic (yet unknown), uncorrelated and slowly varying; under the model (4), noisy estimations of the jitter samples values at pilot positions can be obtained from (11) as $\widetilde{\xi}_{n}=\xi_{n}+w_{n}/x_{n}^{\prime}$. The additive noise is conditionally Gaussian but heteroscedastic because its variance is time-dependent; moreover the (unknown) values $x^{\prime}_{n}$ are to be replaced with their estimations $y^{\prime}_{n}$. We set $v_{n}=w_{n}/y^{\prime}_{n}$. The idea is to perform an optimal short-time ($=$ blockwise) polynomial interpolation of the jitter process at query points $\widetilde{\xi}_{n}$, $n\in I$ for a certain $I$.
Let $N$ be the total number of samples observed; since we want the pilot samples to be evenly spaced, we fix a value $K\in\mathbb{N}$ (the spacing between pilot samples) and set
$I=\{nK+1\,:\,n=0,\dots,\lfloor(N-1)/K\rfloor\};$
we can write $I=\bigcup_{l=1}^{L}I_{l}$, where each $I_{l}$ is a (sorted) set of increasing integer indices so that the last element of $I_{l}$ coincides with the first of $I_{l+1}$. Moreover we require that each $I_{l}$ contains the same number $C$ of elements, except at most for $I_{L}$. We will indicate indices belonging to $I_{l}$ with $i_{c}^{(l)}\in I_{l}$. As mentioned earlier in this very section, we make here the assumption that there exists $L$ vectors $\overline{\boldsymbol{\beta}}^{(l)}$ of polynomial coefficients such that $\mathbf{M}\overline{\boldsymbol{\beta}}^{(l)}=\boldsymbol{\xi}^{(l)}$, where $\mathbf{M}$ is the rectangular measurements matrix
$\mathbf{M}=\begin{bmatrix}1&i_{1}^{(1)}T_{s}&\dots&(i_{1}^{(1)}T_{s})^{d}\\ 1&i_{2}^{(1)}T_{s}&\dots&(i_{2}^{(1)}T_{s})^{d}\\ \vdots&\vdots&\ddots&\vdots\\ 1&i_{C}^{(1)}T_{s}&\dots&(i_{C}^{(1)}T_{s})^{d}\end{bmatrix}$ (12)
and $\boldsymbol{\xi}^{(l)}=(\xi_{i_{1}^{(l)}},\dots,\xi_{i_{C}^{(l)}})^{t}$. Our cost function is the conditional expected error
$\mathbb{E}[\|\mathbf{M}\mathbf{S}\widetilde{\boldsymbol{\xi}}^{(l)}-\boldsymbol{\xi}^{(l)}\|_{2}^{2}\,|\,\widetilde{\mathbf{y}}]$ (13)
as function of $\mathbf{S}$, that we seek to minimize. We cannot use the ordinary least-squares because the noise in the measurements $\widetilde{\boldsymbol{\xi}}^{(l)}$ has time-varying variance. Equation (13) can be written as standard textbook bias + variance decomposition (by expanding the square)
$\|(\mathbf{I}-\mathbf{M}\mathbf{S})\boldsymbol{\xi}^{(l)}\|_{2}^{2}+\mathbb{E}[\|\mathbf{M}\mathbf{S}\mathbf{v}^{(l)}\|_{2}^{2}\,|\,\widetilde{\mathbf{y}}]$ (14)
and by imposing the (necessary) unbiasedness condition $\mathbf{S}\mathbf{M}=\mathbf{I}$ (cfr. (3.28) in [33]) we find that the second term in (14) is minimized if
$\widehat{\mathbf{S}}_{l}=(\mathbf{M}^{t}\mathbf{W}_{l}\mathbf{M})^{-1}\mathbf{M}^{t}\mathbf{W}_{l}$ (15)
where $\mathbb{E}[\mathbf{v}^{(l)}(\mathbf{v}^{(l)})^{t}]=\mathbf{W}_{l}\propto\operatorname{diag}\bigl{\{}(y^{\prime}_{i^{(l)}_{c}})^{2}\bigr{\}}_{c=1}^{C}$, while the bias term vanishes. Notice that $\widehat{\mathbf{S}}_{l}\mathbf{M}=\mathbf{I}$; in addition we recall that (15) solves the weighted least squares problem for heteroscedastic noise [1].
The sought polynomial coefficients are thus computed via $\widehat{\boldsymbol{\beta}}^{(l)}=\widehat{\mathbf{S}}\widetilde{\boldsymbol{\xi}}^{(l)}$, leading to an explicit formula for the interpolating / denoising polynomial functions:
$p_{l}(t)\coloneqq\widehat{\beta}^{(l)}_{1}+\widehat{\beta}^{(l)}_{2}t+\dots+\widehat{\beta}^{(l)}_{d+1}t^{d}.$
The estimated jitter process at all time instants will hence be
$\widehat{\xi}_{n}=\begin{cases}p_{1}(nT_{s}),&1\leq n\leq i_{C}^{(1)}\\ p_{l}((n-i_{1}^{(l)}+1)T_{s}),&i_{1}^{(l)}<n\leq i_{C}^{(l)},\ l=2,\dots,L.\end{cases}$ (16)
In conclusion the de-jittered signal can finally be estimated via:
$\widehat{x}_{n}=y_{n}-\widehat{\xi}_{n}y^{\prime}_{n}.$ (17)
We summarize the algorithm in the following flowchart:
Algorithm 1 Robust polynomial interpolation
1:Observations $y_{n}$, $y^{\prime}_{n}$, pilot indices $I$ and $x_{n}$ for $n\in I$; number of samples $N$, measurements matrix $\mathbf{M}$ (12).
2:
3:for $l=1$ to $L$ do
4:    Form measurements vector $\widetilde{\boldsymbol{\xi}}^{(l)}$ according to (11)
5:    Minimize (13) in $\mathbf{S}$: $\widehat{\mathbf{S}}_{l}\leftarrow(\mathbf{M}^{t}\mathbf{W}_{l}\mathbf{M})^{-1}\mathbf{M}^{t}\mathbf{W}_{l}$
6:    Interpolating poly coefficients $\widehat{\boldsymbol{\beta}}^{(l)}\leftarrow\widehat{\mathbf{S}}\widetilde{\boldsymbol{\xi}}^{(l)}$
7:    Interpolated values at non pilots pos. $\widehat{\boldsymbol{\xi}}^{(l)}$ (16)
8:end for
9:
10:Return de-jittered signal $\widehat{x}_{n}=y_{n}-\widehat{\xi}_{n}y^{\prime}_{n}$ (17) $\forall\,n$.
III-B Kalman smoother
We apply in this section Kalman filtering and smoothing to our jitter tracking problem, that are standard and widely used techniques offering optimal recursive estimations for systems described using the state-space model formalism. The state-space model for the linearized jitter tracking problem under investigation is
$\begin{cases}\xi_{n}=\varphi\xi_{n-1}+\epsilon_{n}\\ y_{n}=x_{n}+\xi_{n}x^{\prime}_{n}+w_{n},\end{cases}$ (18)
and it is readily seen how Kalman filtering and smoothing can estimate the process $\xi_{n}$ as long as a good (to be clarified in which sense) approximation of the parameters $\varphi$, $\sigma_{\epsilon}^{2}$ and white noise variance $\sigma_{w}^{2}$ are available.
Algorithm 2 (Inexact) Kalman Filter and Smoother flow for (18)
1:Observations $y_{n}$, $y^{\prime}_{n}$, pilot indices $I$ and $x_{n}$ for $n\in I$; number of samples $N$; AR(1) parameter $\varphi$; AR proc. variance $Q_{n}=\sigma_{\epsilon}^{2}$; meas. noise variance $R_{n}=\sigma_{w}^{2}$ for $n\in I$, $R_{n}=+\infty$ elsewhere; initial state $\widehat{\xi}_{0|0}$ and covariance $P_{0|0}$.
2:
3:Forward Pass (Inexact Kalman Filter)
4:for $n=1$ to $N$ do
5:    if $n>1$ then
6:        Prediction:
7:        $\widehat{\xi}_{n|n-1}\leftarrow\varphi\widehat{\xi}_{n-1|n-1}$
8:        $P_{n|n-1}\leftarrow\varphi^{2}P_{n-1|n-1}+Q_{n}$
9:    else
10:        Set $\widehat{\xi}_{n|n-1}\leftarrow\widehat{\xi}_{0|0}$
11:        Set $P_{n|n-1}\leftarrow P_{0|0}$
12:    end if
13:    Measurement Update:
14:    Define $H_{n}\leftarrow y_{n}^{\prime}$
15:    Innovation: $\widetilde{y}_{n}\leftarrow y_{n}-x_{n}-H_{n}\widehat{\xi}_{n|n-1}$
16:    Innovation covariance: $S_{n}\leftarrow H_{n}^{2}P_{n|n-1}+R_{n}$
17:    Kalman gain: $K_{n}\leftarrow P_{n|n-1}H_{n}/S_{n}$
18:    Update state estimate: $\widehat{\xi}_{n|n}\leftarrow\widehat{\xi}_{n|n-1}+K_{n}\widetilde{y}_{n}$
19:    Update error covariance: $P_{n|n}\leftarrow P_{n|n-1}-K_{n}H_{n}P_{n|n-1}$
20:end for
21:
22:Backward Pass (Kalman Smoother)
23:Let $\widehat{\xi}_{N|N}$ be the final filter estimate from above
24:$\widehat{\xi}_{N|N}^{\text{smooth}}\leftarrow\widehat{\xi}_{N|N}$
25:for $n=N-1$ down to $1$ do
26:    Compute smoother gain: $C_{n}\leftarrow\varphi P_{n|n}/P_{n+1|n}$
27:    Smoothed state: $\widehat{\xi}_{n|N}^{\text{smooth}}\leftarrow\widehat{\xi}_{n|n}+C_{n}(\widehat{\xi}_{n+1|N}^{\text{smooth}}-\widehat{\xi}_{n+1|n}).$
28:end for
29:
30:Return de-jittered signal (cfr. 17): $\widehat{x}_{n}\leftarrow y_{n}-\widehat{\xi}_{n|N}^{\text{smooth}}y^{\prime}_{n}$
Our implementation of the Kalman filter and smoother routine is described in Algorithm 2. Even though it pretty much follows the standard textbook routine (see e.g. [37]), a slight modification was engineered to handle missing data at non pilots positions: the measurements covariances $R_{n}$ are there set to be =+absent=$\infty$= + ∞, making the Kalman gain null. This correctly encodes the fact that the actual tracked autoregressive process is a subsampled version of the original available through pilots; more concretely the Kalman filter will optimally predict the process $\xi_{n}=\varphi^{K+1}\xi_{n-K-1}+\nu_{n}$ with $\nu_{n}\sim\mathcal{N}(0,(1-\varphi^{2(K+1)})\sigma_{\epsilon}^{2}/(1-\varphi^{2}))$, and the smoother will interpolate between them using the assumed dynamics (18). Notice that the exact observation “matrices” $x_{n}^{\prime}$ are replaced with their estimations $y^{\prime}_{n}$. The flowchart 2 details out and clarifies our implementation.
III-C Complexity analysis
The Kalman smoother processes each sample within the stream; a single iteration in both forward and backward passes has constant cost, thus the total cost for $N$ samples is $O(N)$.
In the polynomial technique, the matrix to be inverted in (15) is $d\times d$; the full cost of (15), including matrix multiplications, is thus $O(Cd^{2}+d^{3})$. Since this is repeated $L$ times, the total cost for calculating the polynomial coefficients will be $O(L(Cd^{2}+d^{3}))$; the polynomial evaluations cost $O(Nd)$, bringing the total complexity to $O(Nd+L(Cd^{2}+d^{3}))$.
IV Simulation results
Figure 4: Varying NDR analysis.
Our test signal $x$ is throughout most of this section is a bandlimited Gaussian process (cfr. Section II) with cutoff frequency $40$ MHz, then sampled at $100$ MS/s and with normalized power so that $\sigma_{x}^{2}=1$. The digital signal is generated by applying the desired brickwall filter to a sequence $\mathbf{x}$ of $N=2^{18}$ independent real-valued Gaussian random variables; the samples of $\mathbf{x}$ are assumed to be taken at the time instants $nT_{s}$ and the jittered samples $x(nT_{s}+\xi_{n})$ can be obtained by processing the analog counterpart of $\mathbf{x}$ via the standard Whittaker-Shannon interpolation formula so that
$x(nT_{s}+\xi_{n})=\sum_{m=-\infty}^{\infty}x_{m}\operatorname*{sinc}\left(\frac{nT_{s}+\xi_{n}-mT_{s}}{T_{s}}\right).$ (19)
Moreover the sequence of known pilot samples $x_{n}$ for $n\in I$ is time-multiplexed with the received signal samples. For the experiments within this section we chose a constant sequence of pilot samples that left unchanged the signal’s variance. In a first batch of simulations we test the interplay between background white noise and jitter, and its effects on the proposed algorithms. In Figure 4 we keep constant (to two different levels) the amount of jitter. We sweep through $30$ different white noise power levels (and run $5$ experiments each) so that the NDR is between $-20$ and $10$ dB and the overall SINADR (cfr. formula (6)) is assessed. Two different pilot samples densities are considered too, while keeping fixed both the blocks length $C=500$ and the polynomial degree $d=4$ for Algorithm 1. The Kalman smoother has oracle knowledge of $\sigma_{w}^{2}$, $\sigma_{\epsilon}^{2}$ and $\varphi=0.999$. As expected, the Kalman smoother displays superior performances throughout the entire NDR range; improvements after compensation seem to be achievable up to $1-2$ dB, after which the jitter disturbance is essentially drowned in noise and thus no longer distinguishable (and leading).
[…]
Another batch of simulations is showed in Figure 5 where now it is the NDR to be kept constant, while the density of the (uniformly spaced) pilot samples is gradually increased from $1\%$ to $20\%$. Both techniques increasingly benefit from the higher number of available pilots, with a remarkable jump in performance for Algorithm 1 when going from $1$ to $2\%$; at the same time the Kalman smoother displays strong estimation capabilities even with pilot density as low as $1\%$.
[…]
In Figure 6 we show the algorithms performances with increasing levels of jitter and constant noise power; the latter is fixed so that the NDR is $-10$ dB at the lowest jitter level. Compensated signals show $6-15$ dB improvements in SINADR up to $\approx 4\%$ jitter, where the Kalman smoother is taken over by the polynomial method that keeps delivering a consistent improvement up to $10\%$ of jitter. Interestingly, beyond a certain point the Kalman smoother’s performance degrades rapidly. This stems from how it assesses the trustworthiness of measurements via the ratio $\sigma_{\epsilon}^{2}/\sigma_{w}^{2}$, and from the fact that in high‑SNR, low‑NDR scenarios (as in Figure 6), the Kalman gain approaches unity - making the posterior essentially the raw measurement and thus highly sensitive to outliers. Of course white noise is one reason for outliers, but when the jitter is large and the NDR small, there are other distortion sources that the Kalman scheme is not actively accounting for, most prominently the approximation $x^{\prime}_{n}\approx y^{\prime}_{n}$ and the increasingly weakening of the linear approximation model (4).
We note that the parameters $d$, $C$ and the pilot density interact in the polynomial technique. In these experiments, we did not attempt to identify their optimal combination; instead, we chose sub-optimal values empirically to avoid both over-fitting and under-fitting. The problem of optimally interpolate Wiener-type processes with polynomial functions has been already studied in e.g. [6]. Thus a more rigorous investigation of this interplay is left to future work.
A last numerical simulation is shown in Figure 7, where we sweep through different signal bandwidths to show the SINADR of the compensated signals, in a similar fashion as in Figure 4; as the bandwidth increases we kept either the NDR (left) or the SNR (right) constant.
Figure 7: Algorithm performances vs signal bandwidth. Fixed NDR $=-5$ dB (left) and fixed SNR $=40$ dB (right). $\varphi=0.999$, $\sigma_{\xi}/T_{s}=1.5\cdot 10^{-2}$.
IV-A Suppression of jitter from blockers
The numerical examples displayed in the previous section deal with jitter suppression seen as in-band distortion; the power spectral densities of compensated and uncompensated signals look mostly the same because the term $\xi x^{\prime}$ is much weaker (in variance) than the signal itself if (II-B) holds; this can be rigorously proved by modifying the proof of Proposition II.2. The fact that the term $\xi x^{\prime}$ produces mostly in-band distortions is not entirely obvious, because its PSD is not compactly supported. However its out-of-band contributions are negligible; this can be seen by studying the integral function
[…]
that has two peaks at $|\nu|\approx\overline{\omega}$ and sharply declines to 00 as soon as $|\nu|>\overline{\omega}$.
To be better able to visually show the spectral impact of the proposed methods, we outline a different (and somewhat, in this section, ideal) scenario borrowed from radar applications; we assume a bistatic pair consisting of an illuminator of opportunity (IO) at known location and an antenna receiver with two channels, one for surveillance (SC) and one for reference (RC). The RC is responsible for acquiring high fidelity “copies” of the IO waveform, while the SC is primarily responsible for capturing targets’ echo. The SC is however also corrupted by Direct-Path Interferences (DPI) which are usually much stronger than echoes, and can thus entirely overshadow them.
The signal after ADC sampling can be written as
$z_{n}=x_{n}+\xi_{n}x_{n}^{\prime}+\zeta_{n}+\xi_{n}\zeta_{n}^{\prime}+w_{n}$ (21)
where $\zeta_{n}$ is the contribution from the DPI while $x_{n}$ is the wanted signal, and $\zeta_{n}+x_{n}$ are assumed to be known if $n\in I$; by hypothesis $\sigma_{\zeta}^{2}\gg\sigma^{2}_{x}$ and $\sigma_{\zeta^{\prime}\xi}^{2}>\sigma^{2}_{x}$ so that even the jitter coming from the blocker overpowers the signal of interest.
Figure 8: PSD at different compensation stages.
We picture a situation in which the term $\zeta_{n}$ can be entirely canceled with negligible impact on the others (at least if the target is moving and there is a nontrivial doppler component in $x_{n}$), for example projecting it away using subspace methods in combination with the RC signal knowledge, see e.g. [10]. After de-jittering $z_{n}$ using the methods presented in the previous section and optimally interpolating the “holes” left after removing the pilot samples using the Gerchberg-Papoulis algorithm, we then project away $\zeta_{n}$ and what is left is the sought echo signal. In Figure 8 are shown examples of power spectral densities of the different compensation stages described in this section.
V Conclusions and future research
In the present paper we proposed two algorithms to address the stochastic distortion appearing in baseband signals produced by local oscillators in Analog-to-Digital Converters. A mathematically clean and convenient description of the jitter noise as autoregressive process of order $1$ was also reiterated. The Kalman smoother approach shows robust and nearly optimal performances in a wide range of scenarios (cfr. Figure 4), and both methods display significant improvements in SINADR levels even with very low sample pilots densities (Figure 5). Their impact is assessed also when pushing the validity of the linear approximation (4) to a breaking point, and we discovered that for very large jitter the polynomial compensation algorithm could be more suitable, as less sensitive to outliers (Figure 6). Future research will focus on assessing the robustness of the Kalman smoother in the face of parameter inaccuracies, on incorporating additional distortions (e.g. phase noise when the signal is upconverted to radio frequency, strong nearby blockers already contaminated with Digital-to-Analog Converter (DAC) jitter, power amplifiers nonlinearities, harmonic distortions etc.) in the compensation routines as well as on designing optimal training sequences of pilot samples.
VI Acknowledgements
ChatGPT “gpt-o4-mini” was used to a limited extent for occasional text harmonization and grammar checks. The authors confirm that all theoretical and numerical contributions, proofs, codes and conclusions are exclusively their own.